# A flexible and general semi-supervised approach to multiple hypothesis testing

### v1

@@ 4.2 RESET controls the FDR or FDP
Let $N$ be the indices of the true null hypotheses. The labels $\{L_{i}:i\in N\}$ are i.i.d ±$\pm 1$ random variables with $\mathbb{P}(L_{i}=1)\leq c_{0}$ independently of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
@@
Assume that the true null p-values are mutually independent and independent of the false null p-values, all the side information, and assume the true null p-values have a non-decreasing density. Then Assumption 3 is satisfied with $c_{0}=\frac{a}{a+b_{2}-b_{1}}$.
@@
In practice, Assumption 3 (or the conditions of Lemma 1) needs to be verified. We provide reasonable justification of this assumption in our real data applications in Sections 5.2 and 6.2. We are now ready to state our two main theorems.
@@
Under Assumption 3 RESET controls the FDR at the user-specified threshold of $\alpha$.
@@
Under Assumption 3 RESET controls the FDP at the user-specified threshold of $\alpha$ with confidence $1-\gamma$.
We defer the main proof of RESET’s FDR and FDP control to Supplementary Section S.3. The proof requires an intermediate Lemma (Lemma S1) which states that the labels $L$ of the true null pseudo targets in Step 4 remaim i.i.d. ±$\pm 1$ RVs independently of everything else, even after sampling the training decoys in Step 2 and the application of the machine learning model in Step 3.
@@ 4.3 Semi-supervised approach
In this section we describe our specific framework for RESET’s Step 3, which as mentioned, can use any semi-supervised machine learning model. It consists of two man steps: (1) we apply a range of classification algorithms to estimate a high quality positive set from the data, containing presumably many false nulls, and (2) this positive set is then subsequently used for a second application of the classification algorithms to rescore the hypotheses.
@@
    Define the negative set as the set of all training decoys and the positive set as a subset of the top scoring pseudo targets (which we elaborate in Step III at the end of this section). We define the features that are subsequently used by a collection of classification algorithms as the combined score and side information $(W,\mathbf{x})$.
@@
    For each considered classification algorithm, we randomly split all the hypotheses into $K$ folds (we used $K=3$ throughout). For each $k\in\{1,2,\dots,K\}$, we train the classification algorithm using the positive and negative set on the features in the $K\setminus\{k\}$ folds and apply the trained model to all the hypotheses in the $k$th test fold. We consider a random forest (RF), a generative additive model (GAM) and a collection of two-layer neural networks (NN) with decay parameter $\lambda\in\{0,0.1,1\}$ and a number of hidden layer nodes $h\in\{2,5,10\}$.
@@
    To evaluate each classification algorithm, we apply Selective SeqStep (SSS) (Algorithm S1 which is the same as SSS+ of Equation (2) without the “+1”) at an FDR threshold $\alpha$ to each of the $rK$ test folds, using the scores obtained from the classification algorithm and the pseudo labels. If $\mathbb{P}(L_{i}=1)\leq c_{0}$ then we set $c=1-s\cdot(1-c_{0})$. This is to ensure that the ratio $\frac{c}{1-c}$ correctly accounts for the number of null pseudo targets to training decoys [^4] This is not a requirement for error control..
@@
    Using the chosen classification algorithm, we assign new scores $\tilde{W}$ to the hypotheses. $\tilde{W}_{i}$ is defined as the average decision value for the $i$th hypothesis, taken over the $K$ test folds that the $i$th hypothesis appears in, and the $r$ repetitions executed in Step iii.
@@
    The positive set is redefined as those pseudo targets that are ‘discovered’ in Step vii. Steps ii-vi are then repeated with the new positive set. To ensure the positive set is not too small, we increase $\alpha$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo targets are in the positive set. We used min_positive = 50.
@@
    To reduce the effect of ‘noisy’ side information that is unable to distinguish between true and false nulls, we implement an initial side information selection procedure. That is, we use a generative additive model to fit a smoothing spline using $W\cdot\tilde{L}$ as the response against each user-provided side information variable one at a time. We use tryCatch in case the smoothing spline fails, in which case we fit the response directly to the side information variable, that is, we use a linear fit. Again, we try to make use of the zero-scoring hypotheses where we can by setting $W\cdot\tilde{L}=0$ in these cases. We keep only the side information variables with sufficiently small p-values in the fitted model ($p<0.01$).
@@
    Not all the pseudo targets need to be included in the initial positive set in Step (i). After all, we expect many of the pseudo targets to be true nulls. Hence, to improve the performance of RESET, we consider the following strategy. For each feature in $(W,\mathbf{x})$, reorder the hypotheses according to that feature, and apply SSS to the pseudo labels $\tilde{L}$ at the FDR level $\alpha$. Then select the feature that maximizes the number of pseudo discoveries. Then using the chosen feature, we define the positive set as those pseudo targets discovered at $\alpha_{0}$ (we used $\alpha_{0}=50\%$ throughout), again using SSS with the pseudo labels $\tilde{L}$. Usually the score is selected, but in some cases the side information is highly informative and is selected instead.
    Users interested in a small FDR level, like $\alpha=1\%$, could probably benefit from reducing $\alpha_{0}$ to further improve the performance, and in particular the speed of RESET (though as we mentioned, we kept $\alpha_{0}=50\%$ across the board). To ensure the positive set is not too small, we increase $\alpha_{0}$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo targets are in the positive set.
We refer to the general RESET wrapper in combination with our semi-supervised approach as RESET Ensemble. We refer to RESET RF/GAM/NN when the ensemble of machine learning methods in Step 2 is reduced to a random forest, generative additive model, and a collection of neural networks as described, respectively.
@@ 4.4 Code availability
Our code is available at https://github.com/freejstone/stat_RESET_paper_code. We intend on developing a package for RESET in R.
@@ 5.1 Numerical experiments
In this section, we conduct a wide range of simulations in the variable selection problem based on those introducted by Ren and Candès [43]. We compare RESET Ensemble to Adaptive Knockoffs (AdaKO) and the generic Knockoff Filter (KO) (see Section 2). Further implementation details for AdaKO are given in Supplementary Secton S.5. We provide a brief description of each simulation setup, referring to [43] for further details. Each setup is designed to satisfy Assumption 3, which is required for our theoretical guarantees. The labels and scores are obtained according to Equation (1). We focus on FDR control in this section, and study FDP control in Section 5.2.
@@ 5.1.1 Simulation 1: Linear model with one-dimensional side information
Our first simulation is based on Simulation 1 by Ren and Candès [43]. They used $n=1000$ independent observations from the joint data distribution of $(X,Y)$, where $X$ is drawn from a hidden markov model (HMM) with $p=900$ variables and $Y|X\sim\mathcal{N}(X\beta,1)$ is a linear model. The $\beta$’s are selected in the following way. Let $k$ denote the number of nonzero $\beta$’s. Then $k$ indices from $\{1,2,\dots,p\}$ are sampled without replacement with indices $i\in\{1,\dots,2k\}$ having probability proportional to $1/i^{2}$ and the indices $i\in\{2k+1,\dots,p\}$ having zero probability. The $k$ chosen indices are the nonzero $\beta$’s and are assigned a value of ±$\pm 3.5/\sqrt{n}$ where the signs are chosen i.i.d uniformly. We consider $k\in\{50,150,300\}$ while the original simulation only considers $k=150$. Each hypothesis is supported by the side information $\mathbf{x}_{i}=i$, which should help in detecting the nonzero $\beta$’s as smaller values of $i$ are more likely to correspond to a nonzero $\beta_{i}$. We used 100 independent runs of the above, where each run consisted of drawing those $n$ observations. The original simulation in [43] sampled the $\beta$’s once and fixed them across the 100 runs. Here we redrew the $\beta$’s for each run so that we may average our results over a range of models, as we noticed significant variability between the draws.
@@ 5.1.2 Simulation 2: Linear model with two-dimensional side information
In a second simulation, we considered a larger linear model where $n=2000$ observations are taken from the joint distribution $(X,Y)$ as described in Section 5.1.1, with $p=1800$. Accordingly, we also adjusted the number of false nulls to $k\in\{150,300,450\}$. Each hypothesis is equipped with the following two-dimensional side information. For the $i$th hypothesis, we draw a pair of observations $\mathbf{x}_{i}$ from a bivariate normal $\mathcal{N}(\boldsymbol{\mu},I)$ with $\boldsymbol{\mu}=(0,0)$ for a true null hypothesis and $\boldsymbol{\mu}=\pm(2,2)$ for a false null hypothesis, where the signs are chosen i.i.d uniformly. We used 20 independent runs of this larger simulation instead of 100 where in each run we draw those $n$ observations, $\beta$ coefficients and side information variables.
@@ 5.1.3 Simulation 3: Logistic model with two-dimensional side information
Next, we consider Simulation 2 by Ren and Candès [43]. They generate $n=1000$ observations with $p=1600$ variables sampled from the joint data distribution of $(X,Y)$ where $X$ is distributed according to a multivariate normal distribution and $Y|X\sim\text{Bernoulli}\left(e^{\beta\cdot X}/\left(1+e^{\beta\cdot X}\right)\right)$ is a logistic model (see [43] for further details). Each hypothesis’ side information is a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from the lattice $[-20,19]\times[-20,19]\subseteq\mathbb{Z}^{2}$. The nonzero $\beta$’s are chosen according to the position of $\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ in the lattice. An image of the nonzero $\beta$’s is given in Figure 1. Each nonzero $\beta$ is set to ±$\pm 25/\sqrt{n}$ where the signs are chosen i.i.d uniformly. We used 98 runs of the above setup (since in 2 runs, AdaKO EM failed with an error), where each run consisted of drawing those $n$ observations. The original simulation in [43] sampled the sign of the $\beta$’s once, fixing them across the 98 runs. Instead, we redraw the sign of the $\beta$’s in each of the 98 runs since we found the results significantly varied with those draws.
@@ 5.1.4 Simulation 4: Large $p$ and $n$ and 3-d side information
Finally, in the last simulation, $n=10K$ observations with $p=10K$ variables and $k=0.15p$ relevant variables are drawn from the joint data distribution $(X,Y)$ exactly as described in Simulation 1. Each variable is equipped with the following three-dimensional side information. The first $k$ variables are randomly assigned a drawn from $i\in\{1,\dots,2k\}$ taken without replacement and having probability proportional to $1/i^{2}$. The remaining $p-k$ variables are uniquely assigned to one of the remaining values from $\{1,2,\dots,p\}$ uniformly at random. We expect that most of the nonzero $\beta$’s to have small values of $i$, but the effect of the random assignment makes the resulting side information variable less informative than in Simulation 1 as depicted in Figure 2. This process is repeated three times, thus associating three side information variables with each hypothesis, which taken together, should be reasonably informative. We conducted 5 independent runs where in each run we drew the $n$ observations, $\beta$ coefficients and side information.
@@ 5.1.5 Results
We first compare RESET Ensemble with Adaptive Knockoff’s default method in R, AdaKO GAM, and the generic Knockoff Filter (KO). In Figure 3, we plot the power of each simulation using these methods as a function of the selected FDR thresholds of $\alpha\in\{0.05,0.1,0.2,0.3\}$. The power is calculated as the proportion of relevant variables correctly discovered, averaged over the number of runs used in each simulation. Similarly, in Supplementary Figure S2, we plot the estimated FDR as the proportion of irrelevant variables in the discovery list, averaged over the runs.
@@
Unsurprisingly, RESET Ensemble uniformly dominates KO, demonstrating it successfully leverages the side information. At the same time, AdaKO GAM mostly improves on KO, with some exceptions.
@@
In Simulation 1, we find that at higher FDR thresholds, RESET Ensemble is the preferred approach in terms of power. On the other hand, AdaKO GAM is more powerful than RESET at lower FDR thresholds, particularly in the case of $k=50$ and $\alpha=5\%$. This is not too surprising given the following two reasons. First, RESET trains its classification algorithms on a relatively small positive set of false nulls. Second, RESET incurs a greater cost in terms of power at lower FDR thresholds if the classification algorithm overits and incorrectly scores a knockoff higher than it should. Indeed, at an FDR threshold of $\alpha=5\%$, each knockoff may cost up to $402/\alpha=$ 40 discoveries.
@@
We find in Simulations 2 and 3 that RESET Ensemble is consistently more powerful than AdaKO GAM; in fact, AdaKO GAM appears to essentially have the same power as KO. The result in Simulation 3 is consistent with the findings of Ren and Candès in their analagous simulation. AdaKO GAM does not implement any smooth terms when the number of side information variables is two or more, and is therefore unable to leverage the side information if the true and false null hypotheses are not “easily” separable. Finally in Simulation 4, we see that RESET Ensemble is generally the most powerful, with essentially the same power as AdaKO GAM at $\alpha=5\%$. We hypothesize that the greater number of variables in this example, particularly relevant variables, appears to overcome the relatively poor performance of RESET Ensemble in Simulation 1 at low FDR thresholds.
@@
Figure 4: Relative power and runtimes for each method. In (A), each boxplot describes the relative powers for each method, computed across the combination of simulations and FDR thresholds in Figure 3. (B) shows the log10 runtimes for each method applied at the 5% FDR level for 5 runs of Simulation 4. For readability, the points are jittered in the horizontal direction.
@@
We compared RESET Ensemble with AdaKO GAM because those are the respective default methods. We complement this comparision with Figure 4A where we summarize the performance of all RESET and Adaptive Knockoff methods. Specifically, for each panel and FDR threshold in Figure 3, we determined the maximum power across all methods and computed the relative power of each method to this maximum. The resulting relative powers are reported as boxplots, giving an indication of the performance of each method across the simulations as a whole. A plot of the actual powers for each of the methods is given in Supplementary Figure S3. While in Supplementary Figure S3 we often find a version of AdaKO that is more powerful, or at least comparable to RESET Ensemble, this varies between the versions. On the other hand, RESET Ensemble’s performance is consistent across these simulations, and overall typically achieves the highest relative power when the results are aggregated. Interestingly, RESET NN seems to be be marginally more powerful than RESET Ensemble. While this may be the case, we still recommend RESET Ensemble as the default approach to alleviate the burden of selecting any singular classification algorithm.
@@
One advantage of RESET is its computational speed. Indeed, our implementation of RESET, as described in Section 4.3, has many steps that can be parallelized: the $r$ evaluations of each classification algorithm on the $K$ folds can all be done in parallel using multiple cores. In contrast, Adaptive Knockoffs must iteratively apply a classifiation algorithm for every nonzero scoring hypothesis in the candidate set $\mathcal{S}$ until the estimated FDR is $\leq\alpha$, or all hypotheses are revealed. Figure 4B demonstrates this advantage and shows the (log10) runtimes of RESET using 10-cores and Adaptive Knockoffs at the 5% FDR level across 5 runs of Simulation 4. We find that each version of RESET is significantly faster than each version of Adaptive Knockoffs. RESET Ensemble is around $10^{3}$ to $10^{4}$ times faster than AdaKO EM, $10^{2}$ to $10^{3}$ times faster than AdaKO RF, and about 3-4 times faster than AdaKO LR and AdaKO GAM. In scenarios where AdaKO EM and AdaKO RF are infeasible, but AdaKO GAM and AdaKO LR are unable to leverage the side information as in Simulations 3 and 4, RESET is arguably more suitable. We investigate such a scenario in the following application.

### v2

@@ 4.2 RESET controls the FDR or FDP
Let $N$ be the indices of the true null hypotheses. The labels $\{L_{i}:i\in N\}$ are i.i.d. ±$\pm 1$ random variables with $\mathbb{P}(L_{i}=1)\leq c_{0}$ and are independent of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
@@
Assume that the true null p-values have a non-decreasing density and that they are mutually independent and that they are independent of the false null p-values as well as of all the side information. Then Assumption 3 is satisfied with $c_{0}=\frac{a}{a+b_{2}-b_{1}}$.
@@
In practice, Assumption 3 (or the conditions of Lemma 1) needs to be verified. We provide reasonable heuristic justification of this assumption in our real data applications in Sections 5.2 and 6.2. We are now ready to state our two main theorems.
@@
Under Assumption 3 RESET controls the FDR at the user-specified threshold $\alpha$.
@@
Under Assumption 3 RESET controls the FDP at the user-specified threshold $\alpha$ and confidence $1-\gamma$.
We defer the proof of RESET’s FDR and FDP control to Supplementary Section S.3. The proof requires an intermediate lemma (Lemma S1), which states that the labels $L$ of the true null pseudo-targets in Step 4 remain i.i.d. ±$\pm 1$ RVs independently of everything else, even after sampling the training decoys in Step 2 and the application of the machine learning model in Step 3.
@@ 4.3 Semi-supervised approach
In this section we describe our specific framework for RESET’s Step 3, which as mentioned, can use any semi-supervised machine learning model. It consists of two main steps: (1) we apply a range of classification algorithms to estimate a high quality positive set from the data, containing presumably many false nulls, and (2) this positive set is then subsequently used for a second application of the classification algorithms to rescore the hypotheses.
@@
    Define the negative set as the set of all training decoys and the positive set as a subset of the top scoring pseudo-targets (which we elaborate on in Step III at the end of this section). We define the features that are subsequently used by a collection of classification algorithms as the combined score and side information $(W,\mathbf{x})$.
@@
    For each considered classification algorithm, we randomly split all the hypotheses into $K$ folds (we used $K=3$ throughout). For each $k\in\{1,2,\dots,K\}$, we train the classification algorithm using the positive and negative set on the features in the $K\setminus\{k\}$ folds and apply the trained model to all the hypotheses in the $k$th test fold. We consider a random forest (RF), a generalized additive model (GAM), and a collection of two-layer neural networks (NN) with decay parameter $\lambda\in\{0,0.1,1\}$ and a number of hidden layer nodes $h\in\{2,5,10\}$.
@@
    To evaluate each classification algorithm, we apply Selective SeqStep (SSS) (Algorithm S1, which is the similar to SSS+ of Equation (2) without the ‘+1’) at an FDR threshold $\alpha$ to each of the $rK$ test folds, using the scores obtained from the classification algorithm and the pseudo labels. If $\mathbb{P}(L_{i}=1)\leq c_{0}$ then we set $c=1-s\cdot(1-c_{0})$. This is to ensure that the ratio $\frac{c}{1-c}$ correctly accounts for the number of null pseudo-targets to training decoys [^4] This is not a requirement for error control..
@@
    Using the chosen classification algorithm, we assign new scores $\tilde{W}$ to the hypotheses. $\tilde{W}_{i}$ is defined as the average decision value for the $i$th hypothesis in the test fold that it appears in, taken over the $r$ repetitions executed in Step iii.
@@
    The positive set is redefined as those pseudo-targets that are ‘discovered’ in Step vii. Steps ii-vi are then repeated with the new positive set. To ensure that the positive set is not too small, we increase $\alpha$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo-targets are in the positive set. We used min_positive = 50.
@@
    To reduce the effect of ‘noisy’ side information that is unable to distinguish between true and false nulls, we implement an initial side information selection procedure. That is, we use a generalized additive model to fit a smoothing spline using $W\cdot\tilde{L}$ as the response against each user-provided side information variable one at a time. We use tryCatch in case the smoothing spline fails, in which case we fit the response directly to the side information variable, that is, we use a linear fit. Again, we try to make use of some of the zero-scoring hypotheses setting $W\cdot\tilde{L}=0$ in these cases. We keep only the side information variables with sufficiently small p-values in the fitted model ($p<0.01$).
@@
    Not all the pseudo-targets need to be included in the initial positive set in Step i. After all, we expect many of the pseudo-targets to be true nulls. Hence, to improve the performance of RESET, we consider the following strategy. For each feature in $(W,\mathbf{x})$, reorder the hypotheses according to that feature, and apply SSS to the pseudo labels $\tilde{L}$ at the FDR level $\alpha$. Then select the feature that maximizes the number of pseudo discoveries. Then using the chosen feature, we define the positive set as those pseudo-targets discovered at $\alpha_{0}$, again using SSS with the pseudo labels $\tilde{L}$ (we used $\alpha_{0}=50\%$ throughout). Usually the score is selected, but in some cases the side information is highly informative and is selected instead.
    Users interested in a small FDR level, like $\alpha=1\%$, could probably benefit from reducing $\alpha_{0}$ to further improve the performance, and in particular the speed, of RESET (though as we mentioned, we kept $\alpha_{0}=50\%$ across the board). To ensure that the positive set is not too small, we increase $\alpha_{0}$ by increments of 0.01 until the positive set reaches min_positive or until all the pseudo-targets are in the positive set.
We refer to the general RESET wrapper in combination with the above semi-supervised approach as RESET Ensemble. We refer to RESET RF/GAM/NN when the ensemble of machine learning methods in Step ii is reduced to a random forest, generalized additive model, and a collection of neural networks as described, respectively.
@@ 4.4 Code availability
Our code is available at https://github.com/freejstone/stat_RESET_paper_code with an Apache-2.0 license. We intend to develop a package for RESET in R.
@@ 5.1 Numerical experiments
In this section, we conduct a wide range of simulations in the variable selection problem based on those introduced by Ren and Candès [44]. We compare RESET Ensemble to Adaptive Knockoffs (AdaKO) and the generic Knockoff Filter (KO) (see Section 2). Further implementation details for AdaKO are given in Supplementary Section S.5. We provide a brief description of each simulation setup, referring to [44] for further details. Each setup is designed to satisfy Assumption 3, which is required for our theoretical guarantees. The labels and scores are obtained according to Equation (1). We focus on FDR control in this section, and we study FDP control in Section 5.2.
@@ 5.1.1 Simulation 1: Linear model with one-dimensional side information
Our first simulation is based on Simulation 1 by Ren and Candès [44]. They used $n=1000$ independent observations from the joint data distribution of $(X,Y)$, where $X$ is drawn from a hidden markov model (HMM) with $p=900$ variables and $Y|X\sim\mathcal{N}(X\beta,1)$ is a linear model. The $\beta$’s are selected in the following way. Let $k$ denote the number of nonzero $\beta$’s. Then $k$ indices from $\{1,2,\dots,p\}$ are sampled without replacement with indices $i\in\{1,\dots,2k\}$ having probability proportional to $1/i^{2}$ and the indices $i\in\{2k+1,\dots,p\}$ having zero probability. The $k$ chosen indices are the nonzero $\beta$’s and are assigned a value of ±$\pm 3.5/\sqrt{n}$ where the signs are chosen i.i.d. uniformly. We consider $k\in\{50,150,300\}$ while the original simulation only considers $k=150$. Each hypothesis is supported by the side information $\mathbf{x}_{i}=i$, which should help in detecting the nonzero $\beta$’s, because smaller values of $i$ are more likely to correspond to a nonzero $\beta_{i}$. We used 100 independent runs of the above, where each run consisted of drawing those $n$ observations. The original simulation in [44] sampled the $\beta$’s once and fixed them across the 100 runs. Here we redrew the $\beta$’s for each run so that we may average our results over a range of models, as we noticed that all the methods exhibited significant variability between the draws.
@@ 5.1.2 Simulation 2: Linear model with two-dimensional side information
In a second simulation, we considered a larger linear model where $n=2000$ observations are taken from the joint distribution $(X,Y)$ as described in Section 5.1.1, with $p=1800$. Accordingly, we also adjusted the number of false nulls to $k\in\{150,300,450\}$. Note that unlike the previous simulation, it is not relevant how the $k$ variable indices are selected from $\{1,2,\dots,p\}$. Each hypothesis is equipped with the following two-dimensional side information. For the $i$th hypothesis, we draw a pair of observations $\mathbf{x}_{i}$ from a bivariate normal $\mathcal{N}(\boldsymbol{\mu},I)$ with $\boldsymbol{\mu}=(0,0)$ for a true null hypothesis and $\boldsymbol{\mu}=\pm(2,2)$ for a false null hypothesis, where the signs are chosen i.i.d. uniformly. We used 20 independent runs of this larger simulation instead of 100, where in each run we draw those $n$ observations, $\beta$ coefficients, and side information variables.
@@ 5.1.3 Simulation 3: Logistic model with two-dimensional side information
Next, we consider Simulation 2 by Ren and Candès [44]. They generate $n=1000$ observations with $p=1600$ variables sampled from the joint data distribution of $(X,Y)$, where $X$ is distributed according to a multivariate normal distribution and $Y|X\sim\text{Bernoulli}\left(e^{\beta\cdot X}/\left(1+e^{\beta\cdot X}\right)\right)$ is a logistic model (see [44] for further details). Each hypothesis’ side information is a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from the lattice $[-20,19]\times[-20,19]\subseteq\mathbb{Z}^{2}$. The nonzero $\beta$’s are chosen according to the position of $\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ in the lattice. An image of the nonzero $\beta$’s is given in Figure 1. Each nonzero $\beta$ is set to ±$\pm 25/\sqrt{n}$ where the signs are chosen i.i.d. uniformly. We used 98 runs of the above setup (since in 2 runs, AdaKO EM failed with an error), where each run consisted of drawing those $n$ observations. The original simulation in [44] sampled the sign of the $\beta$’s once, fixing them across the 100 runs. Instead, we redraw the sign of the $\beta$’s in each of the 98 runs since we found that the results of all the methods significantly varied with those draws.
@@ 5.1.4 Simulation 4: Large $p$ and $n$ and 3-d side information
Finally, in the last simulation, $n=10K$ observations with $p=10K$ variables and $k=0.15p$ relevant variables are drawn from the joint data distribution $(X,Y)$ exactly as described in Simulation 1. Each variable is equipped with the following three-dimensional side information. The first $k$ variables are randomly assigned a value drawn from $i\in\{1,\dots,2k\}$ taken without replacement and having probability proportional to $1/i^{2}$. The remaining $p-k$ variables are uniquely assigned to one of the remaining values from $\{1,2,\dots,p\}$ uniformly at random. We expect most of the nonzero $\beta$’s to have small values of $i$, but the effect of the random assignment makes the resulting side information variable less informative than in Simulation 1, as depicted in Figure 2. This process is repeated three times, thus associating three side information variables with each hypothesis, which taken together, should be reasonably informative. We conducted 5 independent runs, where in each run we drew the $n$ observations, $\beta$ coefficients, and side information.
@@ 5.1.5 Results
We first compare RESET Ensemble with Adaptive Knockoff’s default method in R, AdaKO GAM, and the generic Knockoff Filter (KO). In Figure 3, for each simulation we plot the power of each method as a function of the selected FDR thresholds of $\alpha\in\{0.05,0.1,0.2,0.3\}$. The power is calculated as the proportion of relevant variables correctly discovered, averaged over the number of runs used in each simulation. Similarly, in Supplementary Figure S2, we plot the estimated FDR: the proportion of irrelevant variables in the discovery list, averaged over the runs.
@@
Unsurprisingly, RESET Ensemble uniformly dominates KO, demonstrating that it successfully leverages the side information. At the same time, AdaKO GAM mostly improves on KO, with some exceptions.
@@
In Simulation 1, we find that at higher FDR thresholds, RESET Ensemble is the preferred approach in terms of power. On the other hand, AdaKO GAM is more powerful than RESET at lower FDR thresholds, particularly in the case of $k=50$ and $\alpha=5\%$. This is not too surprising for two reasons. First, RESET trains its classification algorithms on a relatively small positive set of false nulls. Second, RESET incurs a greater cost in terms of power at lower FDR thresholds if the classification algorithm overfits and incorrectly scores a knockoff higher than it should. Indeed, at an FDR threshold of $\alpha=5\%$, each knockoff may cost up to $402/\alpha=$ 40 discoveries.
@@
We find in Simulations 2 and 3 that RESET Ensemble is consistently more powerful than AdaKO GAM; in fact, AdaKO GAM appears to essentially have the same power as KO. The result in Simulation 3 is consistent with the findings of Ren and Candès in their analogous simulation. Indeed, AdaKO GAM does not implement any smooth terms when the number of side information variables is two or more, and is therefore unable to leverage the side information if the true and false null hypotheses are not ‘easily’ separable. Finally in Simulation 4, we see that RESET Ensemble is generally the most powerful, with essentially the same power as AdaKO GAM at $\alpha=5\%$. We hypothesize that the greater number of variables in this example, particularly relevant variables, appears to overcome the relatively poor performance of RESET Ensemble in Simulation 1 at low FDR thresholds.
@@
Figure 4: Relative power and runtimes for each method. (A) Each boxplot describes the relative powers for each method, computed across the combination of simulations and FDR thresholds in Figure 3. (B) The log10 runtimes of each method applied at the 5% FDR level for 5 runs of Simulation 4. For readability, the points are jittered in the horizontal direction.
@@
We compared RESET Ensemble with AdaKO GAM because those are the respective default methods. We complement this comparison with Figure 4A where we summarize the performance of all RESET and Adaptive Knockoff methods. Specifically, for each panel and FDR threshold in Figure 3, we determined the maximum power across all methods and computed the relative power of each method to this maximum. The resulting relative powers are reported as boxplots, giving an indication of the performance of each method across the simulations as a whole. A plot of the actual powers for each of the methods is given in Supplementary Figure S3. While in Supplementary Figure S3 we often find a version of AdaKO that is more powerful, or at least comparable to RESET Ensemble, this varies between the versions. On the other hand, RESET Ensemble’s performance is consistent across these simulations, and overall typically achieves the highest relative power when the results are aggregated. Interestingly, RESET NN seems to be marginally more powerful than RESET Ensemble. While this may be the case, we still recommend RESET Ensemble as the default approach to alleviate the burden of selecting any single classification algorithm.
@@
One advantage of RESET is its computational speed. Indeed, our implementation of RESET, as described in Section 4.3, has many steps that can be parallelized: the $r$ evaluations of each classification algorithm on the $K$ folds can all be done in parallel using multiple cores. In contrast, Adaptive Knockoffs must iteratively apply a classification algorithm for every nonzero scoring hypothesis in the candidate set $\mathcal{S}$ until the estimated FDR is $\leq\alpha$, or all hypotheses are revealed. Figure 4B demonstrates this advantage showing the (log10) runtimes of RESET (using 10 of a 20-cores CPU) and of Adaptive Knockoffs at the 5% FDR level across 5 runs of Simulation 4. We find that each version of RESET is significantly faster than each version of Adaptive Knockoffs. RESET Ensemble is around $10^{3}$ to $10^{4}$ times faster than AdaKO EM, $10^{2}$ to $10^{3}$ times faster than AdaKO RF, and about 3-4 times faster than AdaKO LR and AdaKO GAM. In scenarios, as in Simulations 3 and 4, where AdaKO EM and AdaKO RF are infeasible, and AdaKO GAM and AdaKO LR are unable to leverage the side information RESET is clearly more suitable. We investigate such a commonly encountered scenario in the following application.